For each of the following situations with regard to owner's contributions and withdrawals of the business, compute the amount of net income or net loss during June 2018. (Use a minus sign or parentheses for a net loss.) 


a. The owner contributed $20,000 to the business and made no withdrawals. 

Net income (loss) = $ _______ 

b. The owner made no contributions. The owner withdrew cash of $8,000. 

Net income (loss) = $ _______ 



May 31, 2018
June 30, 2018
Total Assets$133,000
$202,000
Total Liabilities77,000
122,000
Answer #1

a) Net Income = $4000

b) Net Income = $32000

Explanations:

Owner's Equity as at May 31,2018 = 133000 - 77000 = 56000

Owner's Equity as as at June 30,2018 = 202000 - 122000 = 80000

*Net Income = 80000 - 56000 - 20000 = 4000

*Net Income = 80000 - 56000 + 8000 = 32000